# Enhanced sensitivity of sub-THz thermomechanical bolometers exploiting vibrational nonlinearity

### v2

@@ Abstract
A common approach to detecting weak signals or minute quantities involves leveraging on the localized spectral features of resonant modes, where sharper lines (i.e. high Q-factors) enhance transduction sensitivity. However, maximizing the Q-factor often introduced technical challenges in fabrication and design. In this work, we propose an alternative strategy to achieve sharper spectral features by using interference and nonlinearity, all while maintaining a constant dissipation rate. Using far-infrared thermomechanical detectors as a test case, we demonstrate that signal transduction along an engineered response curve slope effectively reduces the detector’s noise equivalent power (NEP). This method, combined with an optimized absorbing layer, achieves sub-pW NEP for electrical read-out detectors operating in the sub-THz range.
@@ I Introduction
Transducer-based sensing – relying on the conversion of energy from one form (the measured quantity) to another (read-out signal) – strongly benefits from spectrally sharp transfer functions, as large derivatives translates into large responsivities. The usual route for optimizing the transfer functions relies on devising resonant elements with low energy loss rates; this manifests into narrow linewidths (large Q-factors) and leads to large sensitivities, which have been widely and successfully employed among the others in mass [1], aerostatic pressure [2], gas [3], temperature [4], polarization state [5] and refractive index [6] sensing.
Increasing the Q-factors of nano- and micro-metric sized detectors, whose footprint is suitable for integration in electronic systems, represents a significant technological challenge, both in terms of design and fabrication. Notewhorty, an intense research has recently led to ultra-high Q-factors (exceeding one billion) in micromechanical resonators oscillating at frequencies from hundreds of kHz to MHz, obtained through soft clamping and dissipation dilution design as well as careful fabrication [7, 8, 9, 10]. Although impressive, reaching consistently these values in commercially compatible processes remains a significant challenge: device imperfections or intrinsic physical effects often create bottlenecks, ultimately limiting the maximum sensitivity of commercial devices, which routinely have Q-factor ranging around 104-105.
@@
Improvements based on exploiting Fano lineshapes are still restricted by the limited manipulation of the Fano factor, which usually depends on the coupling between a broad and a narrow linewidth cavity system [17]. Moreover, the maximum slope of a Fano response function is still connected to the resonator’s Q-factor. This further limit can be surpassed by taking advantage of the intrinsic device nonlinearities, which can produce massive linewidth deformations (foldover effect [18]) or lead to multistability[19], with internal transitions between stable solutions leading to, in principle, “infinitely sharp” spectral features.
These concepts have been successfully implemented in several systems including superconductor [20] and semiconductor [21, 22] based single photon detectors. They also have focused applications in micromechanical systems, including mass sensors monitoring the shift of nonlinear resonant frequencies [23, 24], “threshold-based” sensors exploiting bifurcation phenomena for mass [25, 26] and gas [27, 28] limit detection, as well as more structured platforms exploiting exceptional points [29].
In this paper, we show how nonlinear transduction sensing can be employed to enhance the characteristics of far-infrared light detectors based on thermo-mechanical bolometers (TMBs). TMBs have recently emerged as a powerful technology which offers broadband detection at room-temperature, with single-pixel operation at video-rate and faster [30, 31, 32, 33, 34] and the possibility of scaling up the system to focal plane arrays for multiplexed imaging applications [35, 36]. The best TMB devices have a noise-equivalent power (NEP) in a range from a few $\mathrm{pW/\sqrt{Hz}}$ level, to less than $\mathrm{100\,\,pW/\sqrt{Hz}}$, already outperforming some of the commercially available technologies [37].
Our approach to TMBs makes use of silicon nitride (Si3N4) trampoline resonators, which have previously shown a NEP of about $\mathrm{100\,\,pW/\sqrt{Hz}}$ at 20 Hz operating speed in a 1 $\times$ 1 mm sided membrane, illuminated with a 140 GHz source and optically read out via self-mixing interferometry [32].
Compared to our previous results, the devices here investigated have a reduced pixel size and employ specific layers for improved absorptance. In addition, addressing the TMB with metallic wires in a magnetic field, we switched to all-electrical probing via inductive read-out [38], which, respect to the cumbersome optical readout, is better suited for integration and parallelization in large-scale arrays.
Finally, we exploited the material Duffing nonlinearity [39], an effect arising when the restoring force no longer follows Hooke’s law at high levels of strain, leading to an anharmonic response. By applying nonlinear transduction schemes, we achieve an ultrasmall NEP of about $\mathrm{0.1\,\,pW/\sqrt{Hz}}$ in a narrow bandwidth, evaluated under an illuminating 140 GHz source and at room temperature. Combined with an operating speed of 40 Hz, our device characteristics and measurement technique challenge the state of the art for room-temperature bolometric detectors in the sub-THz range.
@@ II.1 Device details
The devices were based on Si3N4 trampoline membrane resonators, which in the last decade have been successfully employed for classical sensing [40] and quantum applications [41, 42]. The basic device geometry consisted of single membranes made of a 300 nm thick stoichiometric silicon nitride, with a 100$\times$100 $\mathrm{\mu}$m central plate hanging on a 300$\times$300 $\mathrm{\mu}$m frame through four 12 $\mathrm{\mu m}$ wide tethers. All membranes had Cr/Au metallic contacts running through the tethers. These granted for a practical all-electrical read-out and actuation [43, 38, 44] employing inductive reading or Lorentz force, respectively, enabled by a $250$ mT magnetic field induced by static Nd magnets. A scanning electron micrograph of a typical device is reported in Figure 1 (a) along with a sketch outlining the read-out circuit.
In order to enhance the device absorbance without degrading its mechanical properties, we introduced “ultra-light” bidimensional layers in the center of the membrane. We have explored the implementation of both metallic and carbon-based absorbers [45]. The choice for the metal was a Cr/Au bilayer with an overall thickness of about 8 nanometers. At such a small thickness scale the metal layer will be a non-uniform film composed by adjacent grains, increasing its sheet resistance compared to its bulk metal value and favouring the approach to the limit value of 188 $\Omega$ which would nominally give perfect impedance matching and the theoretical 50% absorption limit for an isolated layer [46]. Another interesting avenue made use of amorphous carbon materials which can be directly grown on silicon nitride via chemical vapour deposition [47, 48], allowing fine adjustments of the layer conductivity through the precise control of the film thickness. The material of our choice, pyrolitic carbon (PyC), has shown an impressive absorption of $\sim$ 43% in the sub-THz range [49]. Therefore, we realized and investigated two different device implementations, with the same nominal geometrical parameters for the trampoline resonator and different absorbing layers, namely a 2/6 nm Cr/Au thick layer (Au device) and a 18 nm pyrolitic carbon thick layer (PyC device), respectively.
@@ II.2 Detector read-out
The devices were mounted on a ceramic piezoeletric stack actuator (PZT) layer which was used to coherently excite the membrane motion (see Figure 1 (a)). The AC drive voltage applied on the actuator was generated by the reference channel of a lock-in amplifier (LIA), whose input channel was connected to the TMB contact to sense the magnetomotive voltage $\mathrm{\Delta V}$, induced by the modulation of the concatenated magnetic flux due to the vibrational displacement. TMB, actuator and static magnets were placed on a custom-made printed circuit board; this was inserted in a vacuum chamber to reduce atmospheric viscous damping. Optical access was enabled through a cyclic olefin copolymer (COC) window. The vacuum chamber was additionally mounted on a planar moving stage ($xy$ plane), used to scan the TMB position for extended imaging. The illumination was done via a continuous wave, 30 mW, 140 GHz source focused on the device with two parabolic mirrors (not shown in the sketch of Figure 1 (a)).
@@
The concept behind the detection mechanism lies in the shift of the resonant frequency of specific mechanical modes due to thermally induced device deformations, namely thermal expansion and tensile stress reduction. The temperature change is conversely caused by the absorption of the electromagnetic radiation which one wants to detect. Sweeping the driving frequency, it is possible to directly acquire the whole mechanical spectrum and evaluate the frequency shift $\Delta f$ of specific features of the “bright” spectrum ($A_{b}$) from the “dark” ($A_{b}$) one (Spectral Scan, see Figure 1 (c)). In our operating conditions the frequency shift scales linearly with the illuminating intensity, granting the use of this operating mode with signals with large dynamic ranges and enabling its use for the direct acquisition of images, as reported, for example, in Figure 1 (b), where the source focused beam profile has been imaged via spatial scan. The linear response allows a direct conversion of the image from frequency shift to impinging power; here this was done by imposing the proper beam normalization to its total power, which has been independently measured through a calibrated Golay cell detector. Note that the spectral scan is a generally slow detection method, often limited by the slow acquisition time with lock-in amplifier due to the long frequency sweeps, especially when multiple devices are simultaneously investigated.
@@
Faster detection protocols with a reduced dynamic range rely instead on operating with a single frequency demodulating tone $f_{D}$ and exploiting the transduction effect in an open-loop configuration, as illustrated in the transduction scheme of Figure 1 (d). In our device, the transduction detection speed can reach video-rate [32], limited by the thermal relaxation through the tethers. This process, with Q-factors for our devices ranging between $10^{3}$ and $10^{5}$, fully dominates the dynamics of the entire TMB upon illumination. In the transduction scheme, the frequency shift is directly converted in a read-out voltage $\Delta V_{LI}$ which depends on the difference between dark and bright spectrum amplitudes at $f_{D}$, $A_{d}(f_{D})$ and $A_{b}(f_{D})$, respectively. For the detection of weak signals, the bright spectrum can be recast as a frequency shift of the dark spectrum by a vanishing $\delta f$, obtaining:
@@
and giving a read-out voltage directly proportional to the first derivative of the spectral amplitude at $f_{D}$. Changing $f_{D}$ allows one to explore different regions of the signal derivative, which can be very large in asymmetric and nonlinear resonances. In particular, our resonance shows both Fano interference and nonlinear hardening due to the Duffing effect, which is known to produce an increase in the steepness of the spectral features. This is shown in Figure 1 (e), which displays typical resonance lineshapes of a PyC device at different driving strengths: one can see that the asymmetric Fano resonance becomes steeper around 525 kHz due to the Duffing effect at an increased driving voltage amplitude.
@@ III Spectral transduction
Figure 2: (a): typical ON/OFF spectra for the PyC device with a 190 mV piezo driving voltage. (b): device responsivity and noise spectral density in the dark condition. (c): normalized derivative of the modulus of the OFF spectrum reported in (a). (d): spectrogram of PyC device NEP. The minimum NEP can be found around 524.8 kHz for a modulation frequency of 40 Hz.
We evaluated the effect of asymmetric lineshapes on the figure of merit of transduction experiments with our TMB devices. Typical amplitude spectra of the PyC device, demodulated with a sweeping mechanical driving voltage of 190 mV and with on/off source respectively are shown in Figure 2.
Leveraging on the intensity map of Figure 1 (b), we estimated the radiation power illuminating the detector. Considering the small size of our “pixel” compared to the diffraction limited area at this wavelength, i.e. $\lambda^{2}/4\approx 1.15\,\mathrm{mm^{2}}$, we considered the latter as the effective detection area [32], obtaining an effective illuminating power of $P_{i}\sim 0.1\,\mathrm{mW}$.
@@
where the static spectral shift per unit power $\Delta f/P_{i}$ can be estimated from the results of Fig 2 (a). The dynamic responsivity $R$ was then evaluated by combining the previous results with the Bode frequency response, obtained using square wave modulation of the source in a frequency range $f_{M}$ from 1 to 120 Hz, with an external electrical signal provided by the LIA. A typical dynamic responsivity measured at a driving frequency of 524.8 kHz is reported in the left axis of Figure 2 (b). From the Bode plot we can extract a cut-off frequency of about 20 Hz, defined as the -3 dB drop of the signal amplitude, indicated by a red dot in Figure 2 (b).
Next, the device noise spectral density was measured in dark conditions, in the same range of the Bode plot, using the spectral analyzer tool within the lock-in amplifier (UHF 600 MHz Zurich Instruments). A typical measurement is reported on the right axis of Figure 2 (b). The need to operate around the mechanical frequency imposes the injection of a frequency tone to the oscillator; its phase noise leakage represents the main component of the noise spectral density. Its slope of $\propto f^{-3}$ confirms the dominant flicker noise component, as expected from the Leeson’s equation [50]. Other noise contributions can likely originate from the crosstalk in the device electrical connections. Finally, the ratio of noise spectral density to the dynamic responsivity gives an estimate of the Noise-Equivalent power, which in our case is a spectral function of both the modulating frequency $f_{M}$ and driving frequency $f_{D}$, which selects the resonance spectral feature used for transduction.
@@
The full NEP spectrogram for the PyC device excited with a 190 mV driving tone has been reported in Figure 2 (d). It is significant to observe that, at each driving frequency $f_{D}$, the NEP has a minimum around 40 Hz, giving a good operational speed of our device, when compared with a previous report where slightly larger detectors were considered [32]. Additionally, we notice a drastic reduction of the NEP at a driving frequency of $f_{D}$=524.8 kHz, reaching a value below $1\;\mathrm{pW/\sqrt{Hz}}$, significantly reduced with respect to the rest of the spectrum. Unsurprisingly, this $f_{D}$ corresponds to the spectral region with the highest slope, as can be seen from Figure 2 (c), where we reported the normalized modulus of the derivative calculated from the result of Figure 2 (a).
@@
Qualitatively similar effects are present at different driving voltages: higher drives lead to even sharper lineshapes, with a corresponding decrease in the NEP in very narrow spectral regions. Conversely, lower drives produce an increase in the NEP, although with a consistently low value found in a broader spectral range. The latter is linked to the device dynamic range of operation: large spectral regions with significant slopes are most beneficial for using the detector for imaging or as a power meter; at the other limit, with a point-like extremely large slope, we expect the device to operate as a threshold-switching detector, better suited for the recognition of single events (i.e. laser pulses), as will be discussed later.
@@
Figure 3: Normalized vibrational spectrum (a) and its first derivative (b) of the PyC device. Normalized vibrational spectrum (c) and its first derivative (d) of the Au device. The dashed lines indicate in both cases the position of the maximum and minimum derivative, respectively. (e): experimental data points along with interpolated lines illustrating the NEP estimated at the maximum and minimum of the derivative for both devices.
In our operating scheme, the device characteristics can be indeed estimated by directly looking at the derivative of the resonance. Figure 3 (a) shows the normalized mechanical spectrum of the PyC device under a driving voltage of 90 mV; although less pronounced with respect to the data of Figure 2, this excitation configuration still gives a strong asymmetry, spanning from a combination of interference and material nonlinearities. The asymmetry can be made even more manifest by looking at the derivative: the largest negative slope is about 5 times the largest positive one, see the green dots and lines in Figure 3 (b). The Au device shows a qualitatively similar asymmetry, as shown in Figure 3 (c-d), with some differences given by the different experimental assembly resulting in a different Fano factor and a different resonating frequency. For a fair comparison between the devices, we considered the driving strenght leading to the bifurcation point (discussed in more detail later) which, for device made of similar material and resonating at a similar complex frequency, translates into having the same vibrational amplitude. From this normalization we rescaled the driving voltage of the Au device to 40 mV (Figure 3 (c-d)), so that it can be compared to the 90 mV drive of the PyC one (Figure 3 (a-b)). The Au device also exhibits a strong asymmetry (roughly a 1.6 times increase of the negative to the positive slope) and generally larger derivative with respect to the PyC device.
The NEP evaluated by setting $f_{D}$ at the maximum and minimum slopes of both devices is reported in Figure 3 (e) as a function of the modulation frequency of the 140 GHz source. All the NEP curves have a minimum around 40 - 50 Hz, consistent with the results of Figure 2 (d), and show an overall good value which strongly depends of the chosen transducing frequency. In fact, for both devices, the NEP linearly scales with the derivative, as expected from Eq. (2), resulting in a net enhancements of about a factor of 4 and 1.5, respectively, in good agreement with what expected from the derivatives. Note that the enhancements are within the same mechanical spectrum, obtained just by changing $f_{D}$.

### v3

@@ Abstract
A common approach to detecting weak signals or minute quantities involves leveraging the localized spectral features of resonant modes, whose sharper lines (i.e. high Q-factors) enhance transduction sensitivity. However, maximizing the Q-factor often introduces technical challenges in fabrication and design. In this work, we propose an alternative strategy to achieve sharper spectral features by using interference and nonlinearity, all while maintaining a constant dissipation rate. Using far-infrared thermomechanical detectors as a test case, we demonstrate that signal transduction along an engineered response curve slope effectively reduces the detector’s noise equivalent power (NEP), achieving $\mathrm{\sim 30\,pW/\sqrt{Hz}}$ NEP for electrical read-out, sub-THz detectors with an optimized absorbing layer.
@@ I Introduction
Transducer-based sensing – relying on the conversion of energy from one form (the measured quantity) to another (read-out signal) – strongly benefits from spectrally sharp transfer functions, as large derivatives translate into large responsivities. The usual route for optimizing the transfer functions relies on devising resonant elements with low energy loss rates; this manifests into narrow linewidths (large Q-factors) and leads to large sensitivities, which have been widely and successfully employed among the others in mass [1], aerostatic pressure [2], gas [3], temperature [4], polarization state [5] and refractive index [6] sensing.
Increasing the Q-factors of nano- and micro-metric sized detectors, with footprints suitable for integration in electronic systems, represents a significant technological challenge, both in terms of design and fabrication. Noteworthy, an intense research has recently led to ultra-high Q-factors (exceeding one billion) in micromechanical resonators oscillating at frequencies from hundreds of kHz to MHz, obtained through soft clamping and dissipation dilution design as well as careful fabrication [7, 8, 9, 10]. Although impressive, reaching consistently these values in commercially compatible processes remains a significant challenge: device imperfections or intrinsic physical effects often create bottlenecks, ultimately limiting the maximum sensitivity of commercial devices, which routinely have Q-factors ranging around 104-105.
@@
Improvements based on exploiting Fano lineshapes are still restricted by the limited manipulation of the Fano factor, which usually depends on the coupling between a broad and a narrow linewidth cavity system [17]. Moreover, the maximum slope of a Fano response function is still connected to the resonator’s Q-factor. This further limit can be surpassed by taking advantage of intrinsic device nonlinearities, which can produce massive linewidth deformations (foldover effect [18]) or lead to multistability[19], with internal transitions between stable solutions leading to, in principle, “infinitely sharp” spectral features.
These concepts have been successfully implemented in several systems including superconductors [20] and semiconductors [21, 22] single photon detectors. The same ideas have also found focused applications in micromechanical systems, including mass sensors monitoring the shift of nonlinear resonant frequencies [23, 24], “threshold-based” sensors exploiting bifurcation phenomena for mass [25, 26] and gas [27, 28] limit detection, as well as more structured platforms exploiting exceptional points [29].
@@
In this paper, we show how nonlinear transduction sensing can be employed to enhance the characteristics of far-infrared light detectors based on thermo-mechanical bolometers (TMBs). TMBs have recently emerged as a powerful system which offers broadband detection at room-temperature, with single-pixel operation at video-rate and faster [30, 31, 32, 33, 34] and the possibility of scaling up the system to focal plane arrays for multiplexed imaging applications [35, 36]. The best TMB devices have a noise-equivalent power (NEP) in a range from a few $\mathrm{pW/\sqrt{Hz}}$ level, to some $\mathrm{nW/\sqrt{Hz}}$, in some cases outperforming commercially available technologies [37].
Our approach to TMBs makes use of silicon nitride (Si3N4) trampoline resonators, which have previously shown a NEP of about $\mathrm{100\,\,pW/\sqrt{Hz}}$ at 20 Hz operating speed in a 1$\times$1 mm sided membrane, illuminated with a 140 GHz source and optically read-out via self-mixing interferometry [32].
Compared to our previous results, the devices here investigated have a reduced pixel size and employ specific layers for improved absorptance. In addition, addressing the TMB with metallic wires in a magnetic field, we switched to all-electrical probing via inductive read-out [38], which is better suited for integration and parallelization in large-scale arrays, with the trade-off of having a generically higher noise floor.
By applying nonlinear transduction schemes, based on the material Duffing nonlinearity [39], we achieve a NEP of about $\mathrm{30\,\,pW/\sqrt{Hz}}$, evaluated under an illuminating 140 GHz source and at room temperature. Combined with an operating speed of 20 Hz, our device characteristics and measurement technique challenge the state of the art for room-temperature bolometric detectors in the sub-THz range.
It is generally accepted that the high detector sensitivity comes at a price of reduced dynamic range. Extremely sensitive detectors are only used for extremely weak signals and cannot operate in a wide range of intensities due to saturation or even damage. In our case the degree of nonlinearity controlled through the amplitude of exciting voltage can be used to tune the detector from the regime of reasonable sensitivity combined with high dynamic range (low excitation voltage, linear regime) to the state of a high sensitivity required to detect very weak radiation.
@@ II.1 Device details
The devices are based on Si3N4 trampoline membrane resonators, which in the last decade have been successfully employed for classical sensing [40] and quantum applications [41, 42]. The basic device geometry consists of single membranes made of a 300 nm thick stoichiometric silicon nitride, with a 100$\times$100 $\mathrm{\mu}$m central plate hanging on a 300$\times$300 $\mathrm{\mu}$m frame through four 12 $\mathrm{\mu m}$ wide tethers. All membranes have Cr/Au metallic contacts running through the tethers. These grant a practical all-electrical read-out and actuation [43, 38, 44], enabled by a $250$ mT magnetic field induced by static Nd magnets and leveraging inductive reading or Lorentz force, respectively. A scanning electron micrograph of a typical device is reported in Figure 1 (a) along with a sketch outlining the read-out circuit.
In order to enhance the device absorbance without degrading its mechanical properties, we introduced “ultra-light” bidimensional layers in the center of the membrane. We have explored the implementation of both metallic and carbon-based absorbers [45]. The choice for metal was a Cr/Au bilayer with an overall thickness of about 8 nanometers. At such a small scale the metal layer will be a non-uniform film composed by adjacent grains, increasing its sheet resistance compared to its bulk metal value. This favored approaching the limit value of 188 $\Omega$ sheet resistance which would nominally give perfect impedance matching and the theoretical 50% absorption limit for an isolated layer [46]. The other avenue considered makes use of amorphous carbon materials which can be directly grown on silicon nitride via chemical vapour deposition [47, 48], allowing fine adjustments of the layer conductivity through the precise control of the film thickness. The material of our choice, pyrolitic carbon (PyC), has shown an impressive absorption of $\sim$ 43% in the sub-THz range [49]. We realized and investigated two different device implementations, with the same nominal geometrical parameters for the trampoline resonator and different absorbing layers, namely a 2/6 nm Cr/Au thick layer (Au device) and a 18 nm pyrolitic carbon thick layer (PyC device), respectively.
@@ II.2 Detector read-out
The devices were mounted on a ceramic piezoeletric stack actuator layer which was used to coherently excite the membrane motion (see Figure 1 (a)). The AC drive voltage applied on the actuator was generated by the reference channel of a lock-in amplifier (LIA), whose input channel was connected to the TMB contact to sense the magnetomotive voltage $\mathrm{\Delta V}$, induced by the modulation of the concatenated magnetic flux due to the vibrational displacement. TMB, actuator and static magnets were placed on a custom-made printed circuit board; this was inserted in a vacuum chamber to reduce atmospheric viscous damping. Optical access was enabled through a cyclic-olefin-copolymer (COC) window. The vacuum chamber was additionally mounted on a planar moving stage ($xy$ plane), used to scan the TMB position for extended imaging. The illumination was done via a continuous wave, 30 mW, 140 GHz source focused on the device with two parabolic mirrors (not shown in the sketch of Figure 1 (a)).
@@
The concept behind the detection mechanism lies in the shift of the resonant frequency of specific mechanical modes due to thermally induced device deformations, namely thermal expansion and tensile stress reduction. The temperature change is accordingly caused by the absorption of the electromagnetic radiation which one wants to detect. Sweeping the driving frequency, it is possible to directly acquire the whole mechanical spectrum and evaluate the frequency shift $\Delta f$ of specific features of the “bright” spectrum ($A_{b}$) from the “dark” ($A_{d}$) one (Spectral Scan, see Figure 1 (c)). In our operating conditions, the frequency shift scales linearly with the illuminating intensity, allowing the use of this operating mode with large dynamic ranged signals as well as for the direct acquisition of images, as reported, for example, in Figure 1 (b), where the source focused beam profile has been imaged via spatial scan. The linear response allows a direct conversion of the image from frequency shift to impinging power; here this was done by imposing the proper beam normalization to its total power, which has been independently measured through a calibrated Golay cell detector, taking also into account the absorption of the COC window, which at this frequency stands around $50\%$.
Note that the spectral scan is generally a slow detection method, often limited by the acquisition time of lock-in amplifiers due to the long frequency sweeps, especially when multiple devices are simultaneously investigated.
@@
Faster detection protocols with a reduced dynamic range rely instead on operating with a single frequency driving/demodulating tone $f_{D}$ and exploiting the transduction effect in an open-loop configuration, as illustrated in the transduction scheme of Figure 1 (d). In our device, the transduction detection speed can reach video-rate [32], limited by the thermal relaxation time through the tethers. This process, with Q-factors for our devices ranging between $10^{3}$ and $10^{5}$, fully dominates the dynamics of the entire TMB upon illumination. In the transduction scheme, the frequency shift is directly converted into a read-out voltage $\Delta V_{LI}$ which depends on the difference between dark and bright spectrum amplitudes at $f_{D}$, $A_{d}(f_{D})$ and $A_{b}(f_{D})$, respectively. For the detection of weak signals, the bright spectrum can be recast as a frequency shift of the dark spectrum by a vanishing $\delta f$, obtaining:
@@
and giving a read-out voltage directly proportional to the first derivative of the spectral amplitude at $f_{D}$. Changing $f_{D}$ allows one to explore different regions of the signal derivative, which can be very large in asymmetric and nonlinear resonances. In particular, our resonance shows both Fano interference and nonlinear hardening due to the Duffing effect, which is known to produce an increase in the steepness of the spectral features. This is illustrated in Figure 1 (e), which displays typical resonance lineshapes of a PyC device at different driving strengths: one can see that the already asymmetric Fano resonance becomes steeper around 525 kHz due to the Duffing effect for the larger driving voltage amplitude.
@@ III Spectral transduction
Figure 2: (a): typical ON/OFF spectra for the PyC device with a 190 mV piezo driving voltage. (b): dynamic responsivity for the PyC device. The red dot indicates the cut-off frequency. (c): Allan deviation for the PyC device with a demodulation bandwidth of $200\,\mathrm{Hz}$. (d): spectrogram of PyC device NEP. The minimum NEP can be found around 524.26 kHz. (e): normalized derivative and corresponding OFF spectrum (dashed) for the PyC device with a 90 mV piezo driving voltage.
We evaluated the effect of asymmetric lineshapes on the figure of merit of transduction experiments with our TMB devices. Typical amplitude spectra of the PyC device, demodulated with a sweeping mechanical driving voltage of 190 mV and with on/off source, respectively, are shown in Figure 2.
Comparing dark and bright spectrum we can extract the device static responsivity, using the calibrated intensity map of Figure 1 (b) for estimating the radiation power illuminating the detector. Although one could consider the power deposited on the whole membrane area (300 $\mathrm{\mu m}\times$300 $\mathrm{\mu m}$) or the diffraction-limited area $\lambda^{2}/4\approx 1.15\,\mathrm{mm^{2}}$, for the sake of comparison we decided to adopt the same methodology commonly used in the literature [50] and considered the power illuminating the absorbing layer region ($60\,\mathrm{\mu m}$ x $85\,\mathrm{\mu m}$). This yields an illuminating power of $P_{i}\sim 0.26\,\mathrm{\mu W}$ and $\sim 0.48\,\mathrm{\mu W}$ for the PyC and Au device, respectively. The different illumination power is due to the slightly different $xy$ positions where we placed the TMBs.
@@
where the static spectral shift per unit power $\Delta f/P_{i}$ can be estimated from the results of Fig 2 (a). The dynamic responsivity $R$ was then evaluated by combining the previous results with the Bode frequency response, obtained using square wave modulation of the source in a frequency range $f_{M}$ from 1 to 120 Hz, with an external electrical signal provided by the LIA. A typical dynamic responsivity measured at a driving frequency of 524.8 kHz is reported in Figure 2 (b). As expected, the thermal response of the membrane acts as a low-pass filter, with cut-off frequency given by the inverse of thermal relaxation time $\tau_{th}$. The cut-off frequency extracted from the Bode is about 20 Hz, as indicated by a red dot in Figure 2 (b), and corresponds to a $\tau_{th}\sim 50$ ms, compatible with other reports in the literature [36, 50].
Next, the device noise was measured in dark conditions, by calculating the Allan deviation $\sigma_{AD}$, which provides a time-domain characterization of the resonator frequency stability. The Allan deviation was computed from signal recordings over one minute at various driving frequencies, using the UHF Lock-in amplifier (Zurich Instruments), with demodulation bandwidth of $10\,\mathrm{Hz}$, which lies below the thermal cut-off frequency of the device. A representative measurement from a longer acquisition is shown in Figure 2(c), using a demodulation bandwidth of $200\,\mathrm{Hz}$ to show the behavior at lower $\tau$. The typical behavior is observed: at short averaging times $\tau$, the Allan deviation follows a $\propto\tau^{-1/2}$ trend, indicative of white frequency noise dominance, while at longer $\tau$, the impact of frequency drift becomes evident.
Allan deviation and dynamic responsivity can be combined to yield the Noise-Equivalent Power; considering the time-dependence of both the Bode plot of Fig. 2 (b) and the noise contribution of Fig. 2 (c), as expected, the NEP is a function of the frequency of modulation of the incoming signal ($f_{M}$). Moreover, given the strong asymmetry of our lineshapes, in our device the NEP is also function of the mechanical driving frequency $f_{D}$:
$NEP(f_{M},f_{D})=\frac{\sigma_{AD}\sqrt{2\tau}}{R}$ (3)
@@
The full NEP spectrogram for the PyC device excited with a 90 mV driving tone can then be expressed in the $f_{M}-f_{D}$ plane, as reported in Figure 2 (d). It is significant to observe that there is a drastic reduction of the NEP at a driving frequency of $f_{D}$=524.26 kHz, reaching values below $1\;\mathrm{nW/\sqrt{Hz}}$, significantly reduced with respect to the rest of the spectrum. Unsurprisingly, this driving frequency corresponds to the spectral region with the highest slope, as can be seen from Figure 2 (e), where we reported the normalized modulus of the derivative calculated from the resonance lineshape (also reported as a dashed line).
@@
Qualitatively similar effects are present at different driving voltages: higher drives lead to even sharper lineshapes, with a corresponding decrease in the NEP in narrow spectral regions. Conversely, lower drives produce an increase in the NEP, albeit showing the best performances in a broader spectral range. This concept is directly linked to the device dynamic range of operation: large spectral regions with significant slopes are most beneficial for using the detector for imaging or as a power meter; at the other limit, extremely large slope in very narrow frequency ranges, we expect the device to operate as a threshold-switching detector, better suited for the recognition of single events (i.e. laser pulses), as will be discussed later.
@@
Figure 3: Normalized vibrational spectrum (a) and its first derivative (b) for the PyC device. Normalized vibrational spectrum (c) and its first derivative (d) for the Au device. The dashed lines indicate in both cases the position of the maximum and minimum derivative, respectively. (e): NEP evaluated for $f_{D}$ corresponding to the maximum (i.e., most positive) and minimum (i.e., most negative) of the derivative for both devices. The dashed line indicates the theoretical thermal limit of the NEP, calculated for the PyC device.
As we have illustrated, a critical parameter in our operating scheme is the derivative of the mechanical resonance: changing the transduction frequency can dramatically change the device response. For example, Figure 3 (a) shows the normalized mechanical spectrum of PyC device under a driving voltage of 112 mV.
The asymmetry becomes even more manifest by looking at the derivative: the largest negative slope is about 5.6 times the largest positive one, see the green dots and lines in Figure 3 (b). The Au device shows a qualitatively similar asymmetry, as shown in Figure 3 (c-d). The two devices are nominally identical apart from the different absorption layer. Differences in their resonant spectra can be ascribed to the different experimental assembly, starting from the local properties of the piezoelectric actuator where they are glued upon. The piezoelectric stack represents a cross-talk channel strongly impacting the Fano factor, see a more detailed modeling in [36]. For a fair device comparison, we then evaluated the driving voltage leading to the bifurcation point (discussed in more detail later): this acts as a common reference, since we expect both devices to have the same vibrational amplitude at bifurcation. Scaling from this value, we applied a driving voltage of 118 mV to the Au device (Figure 3 (c-d)), so that it can be compared to the 112 mV drive of the PyC one (Figure 3 (a-b)). The Au device also exhibits a strong asymmetry (roughly a 37.5 times increase of the negative to the positive slope) and generally a comparable but smaller derivative with respect to the PyC device.